Lasser Company plans to produce 18,000 units next period at a denominator activity of 54,000 direct labor-hours. The direct labor wage rate is $12.00 per hour. The company's standards allow 2 yards of direct materials for each unit of product; the material costs $8.60 per yard. The company's budget includes variable manufacturing overhead cost of $2.40 per direct labor-hour and fixed manufacturing overhead of $334,800 per period.

Required:

1. Using 54,000 direct labor-hours as the denominator activity, compute the predetermined overhead rate and break it down into variable and fixed elements.

2. Complete the standard cost card below for one unit of product
Answer #1
1
predetermined overhead rate 8.6 per DLH
Variable element 2.4 per DLH
Fixed element 6.2 per DLH
( fixed = 334800/54000 = $6.2 per DLH)
Required 2
Standard Standard Standard
Qty or price or Cost
Hours Rate
Inputs
Direct materials 2 yards $8.6 per yard 17.2
Direct labor 3 hrs $12 per hr 36
Variable manufacturing overhead 3 hrs $2.4 per DLH 7.2
Fixed manufacturing overhead 3 hrs $6.2 per DLH 18.6
Total standard cost per unit 79

Direct labor hours per unit = 54,000/18,000 = 3